OpenStudy (anonymous):

Simplify the expression 11x − (2x^2 + 8x)

OpenStudy (callisto):

11x − (2x^2 + 8x) = 11x -2x^2 - 8x = -2x^2 +11x-8x =?

OpenStudy (anonymous):

this question keeps popping up. yesterday and today remove the parentheses by distributing the minus sign combine like terms

OpenStudy (anonymous):

what callisto said

OpenStudy (anonymous):

-2x2 + 3x????????

OpenStudy (callisto):

Yes, should be

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome